# Are foundation models for computer vision good conformal predictors?

### v2

@@ G Few-Shot Adaptation
ImageNet DINOv2-B 1.36 7.50 1.67 0.900 0.900 0.900 0.414 0.476 0.489
VICReg (RN 50x2) 3.08 13.49 3.89 0.900 0.900 0.900 0.442 0.549 0.445
CLIP 3.03 9.50 3.73 0.900 0.901 0.900 0.434 0.556 0.418
ImageNet-V2 DINOv2-B 1.38 9.47 1.74 0.879 0.892 0.881 0.081 0.187 0.068
VICReg (RN 50x2) 3.15 12.54 3.89 0.867 0.871 0.870 0.130 0.175 0.143
CLIP 2.88 6.79 3.58 0.877 0.864 0.882 0.023 0.034 0.029
ImageNet-Sketch DINOv2-B 1.35 10.90 1.91 0.884 0.904 0.893 0.221 0.322 0.275
VICReg (RN 50x2) 3.49 21.98 4.51 0.826 0.879 0.830 0.119 0.223 0.114
CLIP 3.16 12.98 3.99 0.877 0.900 0.878 0.182 0.293 0.185
ImageNet-A DINOv2-B 1.63 25.30 3.55 0.646 0.897 0.734 0.000 0.000 0.000
VICReg (RN 50x2) 8.43 59.83 7.82 0.531 0.879 0.511 0.000 0.000 0.000
CLIP 6.79 39.42 6.45 0.648 0.912 0.624 0.000 0.000 0.000
ImageNet-R DINOv2-B 1.31 7.49 1.94 0.867 0.910 0.879 0.282 0.532 0.408
VICReg (RN 50x2) 5.84 27.36 5.65 0.761 0.897 0.742 0.198 0.429 0.186
CLIP 3.38 11.93 4.13 0.893 0.930 0.890 0.422 0.604 0.386
(a) DINOv2-B, ImageNet-A
(b) DINOv2-B, ImageNet-R
(c) DINOv2-B, ImageNet-Sketch
Figure 16: Domain shift analysis. Distribution of class-conditional coverages for DINOv2-B on ImageNet-A, ImageNet-R, and ImageNet-Sketch: APS (left) and RAPS (right).
(a) VICReg (RN 50x2), ImageNet-A
(b) VICReg (RN 50x2), ImageNet-R
(c) VICReg (RN 50x2), ImageNet-Sketch
Figure 17: Domain shift analysis. Distribution of class-conditional coverages for VICReg (RN 50x2) on ImageNet-A, ImageNet-R, and ImageNet-Sketch: APS (left) and RAPS (right).
(a) CLIP, ImageNet-A
(b) CLIP, ImageNet-R
(c) CLIP, ImageNet-Sketch
Figure 18: Domain shift analysis. Distribution of class-conditional coverages for CLIP on ImageNet-A, ImageNet-R, and ImageNet-Sketch: APS (left) and RAPS (right).
Figure 19: $\text{ViT}_{\text{ImageNet}}$ vs $\text{ViT}_{\text{MetaCLIP}}$. Analyzing the difference in set size for APS between a $\text{ViT}_{\text{MetaCLIP}}$ and $\text{ViT}_{\text{ImageNet}}$. Equal set sizes not shown.
Figure 20: $\text{ViT}_{\text{ImageNet}}$ vs $\text{ViT}_{\text{DINO-S}}$. Analyzing the difference in set size for APS between a $\text{ViT}_{\text{DINO-S}}$ and $\text{ViT}_{\text{ImageNet}}$. Equal set sizes not shown.
@@
ImageNet ZS 2.81 10.07 3.23 0.086 0.069 0.084
ZSLP 2.17 6.60 2.45 0.076 0.060 0.074
CLAP [53] 2.15 6.83 2.45 0.077 0.060 0.074
CoOp [70] 2.44 6.32 2.79 0.080 0.066 0.080
KgCoOp [17] 2.47 7.18 2.84 0.081 0.065 0.081
-V2 ZS 4.64 17.34 5.72 0.128 0.124 0.128
ZSLP 3.89 12.51 4.81 0.126 0.121 0.125
CLAP [53] 3.72 12.72 4.62 0.126 0.122 0.125
CoOp [70] 4.08 11.26 4.87 0.126 0.123 0.126
KgCoOp [17] 4.09 12.65 4.93 0.126 0.123 0.126
-Sketch ZS 14.74 37.18 20.39 0.099 0.090 0.099
ZSLP 16.46 37.27 22.79 0.098 0.091 0.098
CLAP [53] 15.20 34.90 21.29 0.098 0.090 0.098
CoOp [70] 16.06 34.36 22.12 0.098 0.091 0.099
KgCoOp [17] 15.46 34.35 21.74 0.098 0.092 0.099
-A ZS 9.42 16.59 11.28 0.094 0.085 0.095
ZSLP 10.92 18.63 12.47 0.087 0.080 0.088
CLAP [53] 9.36 16.56 11.38 0.087 0.082 0.087
CoOp [70] 9.23 14.65 10.72 0.087 0.084 0.090
KgCoOp [17] 8.66 14.39 10.19 0.093 0.084 0.096
-R ZS 1.92 5.77 2.31 0.059 0.039 0.054
ZSLP 2.07 5.62 2.36 0.058 0.041 0.056
CLAP [53] 1.92 5.46 2.26 0.059 0.039 0.053
CoOp [70] 2.06 5.11 2.40 0.061 0.042 0.063
KgCoOp [17] 1.96 5.32 2.29 0.060 0.041 0.058
Table 9: Set size and CovGap for CLIP (ViT-B backbone) on ImageNet and its variants.
@@
Aircraft ZS 18.7 19.75 20.74 0.138 0.139 0.127
ZSLP 9.73 11.13 11.55 0.082 0.073 0.086
CoOp [70] 9.41 10.62 11.38 0.081 0.076 0.088
KgCoOp [17] 9.46 10.82 11.0 0.084 0.076 0.086
CLAP [53] 9.73 11.0 11.51 0.082 0.073 0.086
Caltech-101 ZS 0.94 1.95 1.20 0.136 0.092 0.092
ZSLP 0.91 1.20 1.05 0.147 0.085 0.085
CoOp [70] 0.92 1.14 1.06 0.143 0.085 0.087
KgCoOp [17] 0.92 1.25 1.08 0.147 0.084 0.085
CLAP [53] 0.91 1.31 1.08 0.144 0.085 0.086
DTD ZS 11.78 13.3 13.66 0.122 0.117 0.121
ZSLP 4.74 6.83 5.67 0.092 0.077 0.091
CoOp [70] 5.68 6.46 7.00 0.099 0.093 0.098
KgCoOp [17] 5.33 6.22 6.33 0.097 0.083 0.096
CLAP [53] 4.75 6.83 5.68 0.094 0.079 0.093
EuroSAT ZS 4.55 4.78 5.06 0.103 0.101 0.102
ZSLP 1.62 2.22 1.99 0.043 0.051 0.042
CoOp [70] 1.82 2.16 1.95 0.043 0.028 0.034
KgCoOp [17] 1.51 1.92 1.76 0.028 0.030 0.032
CLAP [53] 1.73 2.33 2.06 0.043 0.048 0.041
Flowers-102 ZS 9.26 9.52 15.15 0.180 0.175 0.179
ZSLP 1.09 2.36 1.65 0.132 0.089 0.095
CoOp [70] 1.13 1.80 1.51 0.104 0.080 0.083
KgCoOp [17] 1.19 2.04 1.64 0.114 0.082 0.085
CLAP [53] 1.15 2.47 1.70 0.137 0.092 0.097
Food-101 ZS 1.14 1.89 1.45 0.054 0.026 0.031
ZSLP 1.11 1.82 1.42 0.046 0.023 0.026
CoOp [70] 1.15 1.77 1.41 0.050 0.026 0.030
KgCoOp [17] 1.10 1.86 1.43 0.050 0.025 0.027
CLAP [53] 1.10 1.83 1.42 0.047 0.023 0.026
ImageNet ZS 2.81 10.07 3.23 0.086 0.069 0.084
ZSLP 2.48 8.02 2.88 0.078 0.062 0.078
CoOp [70] 2.69 7.33 3.03 0.081 0.067 0.081
KgCoOp [17] 2.65 8.60 3.01 0.082 0.066 0.082
CLAP [53] 2.44 8.05 2.82 0.079 0.062 0.079
Oxford Pets ZS 1.05 1.43 1.33 0.109 0.067 0.072
ZSLP 0.95 1.26 1.19 0.080 0.039 0.038
CoOp [70] 0.95 1.17 1.13 0.084 0.036 0.036
KgCoOp [17] 0.94 1.20 1.15 0.079 0.038 0.037
CLAP [53] 0.95 1.28 1.20 0.084 0.039 0.038
Stanford Cars ZS 2.24 3.09 2.49 0.109 0.080 0.102
ZSLP 1.62 2.39 1.94 0.083 0.061 0.067
CoOp [70] 1.95 2.64 2.21 0.087 0.065 0.078
KgCoOp [17] 1.99 2.77 2.29 0.096 0.069 0.084
CLAP [53] 1.63 2.40 1.96 0.085 0.061 0.067
SUN397 ZS 2.83 5.80 3.21 0.093 0.073 0.093
ZSLP 2.00 3.96 2.29 0.080 0.060 0.072
CoOp [70] 2.35 3.76 2.64 0.082 0.068 0.081
KgCoOp [17] 2.34 4.18 2.60 0.083 0.065 0.082
CLAP [53] 2.01 3.99 2.31 0.080 0.060 0.072
UCF101 ZS 2.89 5.24 3.83 0.124 0.096 0.124
ZSLP 1.58 3.12 1.98 0.112 0.074 0.088
CoOp [70] 1.97 3.12 2.18 0.114 0.077 0.110
KgCoOp [17] 1.74 3.06 2.06 0.114 0.070 0.096
CLAP [53] 1.63 3.25 2.00 0.114 0.076 0.090

### v3

@@ Appendix F Few-Shot Adaptation
ImageNet DINOv2-B 1.36 $\pm$ 0.00 7.50 $\pm$ 0.02 1.67 $\pm$ 0.00 0.900 $\pm$ 0.000 0.900 $\pm$ 0.000 0.900 $\pm$ 0.000 0.414 $\pm$ 0.005 0.476 $\pm$ 0.005 0.489 $\pm$ 0.005
VICReg (RN 50x2) 3.08 $\pm$ 0.01 13.49 $\pm$ 0.02 3.89 $\pm$ 0.01 0.900 $\pm$ 0.000 0.900 $\pm$ 0.000 0.900 $\pm$ 0.000 0.442 $\pm$ 0.004 0.549 $\pm$ 0.005 0.445 $\pm$ 0.005
CLIP (ViT-B) 3.03 $\pm$ 0.00 9.50 $\pm$ 0.02 3.73 $\pm$ 0.00 0.900 $\pm$ 0.000 0.901 $\pm$ 0.000 0.900 $\pm$ 0.000 0.434 $\pm$ 0.005 0.556 $\pm$ 0.004 0.418 $\pm$ 0.006
ImageNet-V2 DINOv2-B 1.38 $\pm$ 0.00 9.47 $\pm$ 0.04 1.74 $\pm$ 0.00 0.879 $\pm$ 0.000 0.892 $\pm$ 0.000 0.881 $\pm$ 0.000 0.081 $\pm$ 0.009 0.187 $\pm$ 0.009 0.068 $\pm$ 0.009
VICReg (RN 50x2) 3.15 $\pm$ 0.00 12.54 $\pm$ 0.04 3.89 $\pm$ 0.00 0.867 $\pm$ 0.000 0.871 $\pm$ 0.000 0.870 $\pm$ 0.000 0.130 $\pm$ 0.008 0.175 $\pm$ 0.009 0.143 $\pm$ 0.009
CLIP (ViT-B) 2.88 $\pm$ 0.00 6.79 $\pm$ 0.02 3.58 $\pm$ 0.00 0.877 $\pm$ 0.000 0.864 $\pm$ 0.000 0.882 $\pm$ 0.000 0.023 $\pm$ 0.005 0.034 $\pm$ 0.007 0.029 $\pm$ 0.005
ImageNet-Sketch DINOv2-B 1.35 $\pm$ 0.09 10.90 $\pm$ 0.11 1.91 $\pm$ 0.31 0.884 $\pm$ 0.001 0.904 $\pm$ 0.001 0.893 $\pm$ 0.001 0.221 $\pm$ 0.008 0.322 $\pm$ 0.007 0.275 $\pm$ 0.007
VICReg (RN 50x2) 3.49 $\pm$ 0.01 21.98 $\pm$ 0.05 4.51 $\pm$ 0.01 0.826 $\pm$ 0.000 0.879 $\pm$ 0.000 0.830 $\pm$ 0.000 0.119 $\pm$ 0.007 0.223 $\pm$ 0.008 0.114 $\pm$ 0.007
CLIP (ViT-B) 3.16 $\pm$ 0.01 12.98 $\pm$ 0.03 3.99 $\pm$ 0.01 0.877 $\pm$ 0.000 0.900 $\pm$ 0.000 0.878 $\pm$ 0.000 0.182 $\pm$ 0.008 0.293 $\pm$ 0.008 0.185 $\pm$ 0.007
ImageNet-A DINOv2-B 1.63 $\pm$ 1.76 25.30 $\pm$ 1.58 3.55 $\pm$ 1.80 0.646 $\pm$ 0.001 0.897 $\pm$ 0.001 0.734 $\pm$ 0.001 0.000 $\pm$ 0.010 0.000 $\pm$ 0.000 0.000 $\pm$ 0.010
VICReg (RN 50x2) 8.43 $\pm$ 0.03 59.83 $\pm$ 0.07 7.82 $\pm$ 0.01 0.531 $\pm$ 0.001 0.879 $\pm$ 0.001 0.511 $\pm$ 0.001 0.000 $\pm$ 0.000 0.000 $\pm$ 0.000 0.000 $\pm$ 0.000
CLIP (ViT-B) 6.79 $\pm$ 0.02 39.42 $\pm$ 0.07 6.45 $\pm$ 0.01 0.648 $\pm$ 0.001 0.912 $\pm$ 0.001 0.624 $\pm$ 0.001 0.000 $\pm$ 0.000 0.000 $\pm$ 0.000 0.000 $\pm$ 0.000
ImageNet-R DINOv2-B 1.31 $\pm$ 0.05 7.49 $\pm$ 0.04 1.94 $\pm$ 0.24 0.867 $\pm$ 0.001 0.910 $\pm$ 0.001 0.879 $\pm$ 0.001 0.282 $\pm$ 0.006 0.532 $\pm$ 0.009 0.408 $\pm$ 0.007
VICReg (RN 50x2) 5.84 $\pm$ 0.02 27.36 $\pm$ 0.04 5.65 $\pm$ 0.01 0.761 $\pm$ 0.001 0.897 $\pm$ 0.000 0.742 $\pm$ 0.001 0.198 $\pm$ 0.007 0.429 $\pm$ 0.008 0.186 $\pm$ 0.007
CLIP (ViT-B) 3.38 $\pm$ 0.01 11.93 $\pm$ 0.03 4.13 $\pm$ 0.00 0.893 $\pm$ 0.000 0.930 $\pm$ 0.000 0.890 $\pm$ 0.000 0.422 $\pm$ 0.008 0.604 $\pm$ 0.007 0.386 $\pm$ 0.009
Table 11: Set size and CovGap for CLIP (ViT-B) on ImageNet and its variants.
@@
ImageNet ZS 2.81 $\pm$ 0.00 10.07 $\pm$ 0.02 3.23 $\pm$ 0.00 0.086 $\pm$ 0.000 0.069 $\pm$ 0.000 0.084 $\pm$ 0.000
ZSLP 2.17 $\pm$ 0.00 6.60 $\pm$ 0.01 2.45 $\pm$ 0.00 0.076 $\pm$ 0.000 0.060 $\pm$ 0.000 0.074 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 2.44 $\pm$ 0.00 6.32 $\pm$ 0.01 2.79 $\pm$ 0.00 0.080 $\pm$ 0.000 0.066 $\pm$ 0.000 0.080 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 2.47 $\pm$ 0.00 7.18 $\pm$ 0.01 2.84 $\pm$ 0.00 0.081 $\pm$ 0.000 0.065 $\pm$ 0.000 0.081 $\pm$ 0.000
CLAP Silva-Rodríguez et al. (2024) 2.15 $\pm$ 0.00 6.83 $\pm$ 0.01 2.45 $\pm$ 0.00 0.077 $\pm$ 0.000 0.060 $\pm$ 0.000 0.074 $\pm$ 0.000
-V2 ZS 4.64 $\pm$ 0.02 17.34 $\pm$ 0.08 5.72 $\pm$ 0.03 0.128 $\pm$ 0.000 0.124 $\pm$ 0.000 0.128 $\pm$ 0.000
ZSLP 3.89 $\pm$ 0.01 12.51 $\pm$ 0.05 4.81 $\pm$ 0.01 0.126 $\pm$ 0.000 0.121 $\pm$ 0.000 0.125 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 4.08 $\pm$ 0.01 11.26 $\pm$ 0.06 4.87 $\pm$ 0.01 0.126 $\pm$ 0.000 0.123 $\pm$ 0.000 0.126 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 4.09 $\pm$ 0.01 12.65 $\pm$ 0.06 4.93 $\pm$ 0.02 0.126 $\pm$ 0.000 0.123 $\pm$ 0.000 0.126 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 3.72 $\pm$ 0.01 12.72 $\pm$ 0.05 4.62 $\pm$ 0.01 0.126 $\pm$ 0.000 0.122 $\pm$ 0.000 0.125 $\pm$ 0.000
-Sketch ZS 14.74 $\pm$ 0.02 37.18 $\pm$ 0.06 20.39 $\pm$ 0.04 0.099 $\pm$ 0.000 0.090 $\pm$ 0.000 0.099 $\pm$ 0.000
ZSLP 16.46 $\pm$ 0.03 37.27 $\pm$ 0.05 22.79 $\pm$ 0.05 0.098 $\pm$ 0.000 0.091 $\pm$ 0.000 0.098 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 16.06 $\pm$ 0.03 34.36 $\pm$ 0.05 22.12 $\pm$ 0.05 0.098 $\pm$ 0.000 0.091 $\pm$ 0.000 0.099 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 15.46 $\pm$ 0.02 34.35 $\pm$ 0.05 21.74 $\pm$ 0.04 0.098 $\pm$ 0.000 0.092 $\pm$ 0.000 0.099 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 15.20 $\pm$ 0.03 34.90 $\pm$ 0.05 21.29 $\pm$ 0.05 0.098 $\pm$ 0.000 0.090 $\pm$ 0.000 0.098 $\pm$ 0.000
-A ZS 9.42 $\pm$ 0.04 16.59 $\pm$ 0.06 11.28 $\pm$ 0.05 0.094 $\pm$ 0.000 0.085 $\pm$ 0.000 0.095 $\pm$ 0.000
ZSLP 10.92 $\pm$ 0.04 18.63 $\pm$ 0.07 12.47 $\pm$ 0.05 0.087 $\pm$ 0.000 0.080 $\pm$ 0.000 0.088 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 9.23 $\pm$ 0.03 14.65 $\pm$ 0.05 10.72 $\pm$ 0.05 0.087 $\pm$ 0.001 0.084 $\pm$ 0.000 0.090 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 8.66 $\pm$ 0.04 14.39 $\pm$ 0.06 10.19 $\pm$ 0.04 0.093 $\pm$ 0.000 0.084 $\pm$ 0.000 0.096 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 9.36 $\pm$ 0.03 16.56 $\pm$ 0.06 11.38 $\pm$ 0.04 0.087 $\pm$ 0.000 0.082 $\pm$ 0.000 0.087 $\pm$ 0.000
-R ZS 1.92 $\pm$ 0.00 5.77 $\pm$ 0.01 2.31 $\pm$ 0.00 0.059 $\pm$ 0.000 0.039 $\pm$ 0.000 0.054 $\pm$ 0.000
ZSLP 2.07 $\pm$ 0.00 5.62 $\pm$ 0.01 2.36 $\pm$ 0.00 0.058 $\pm$ 0.000 0.041 $\pm$ 0.000 0.056 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 2.06 $\pm$ 0.00 5.11 $\pm$ 0.01 2.40 $\pm$ 0.00 0.061 $\pm$ 0.000 0.042 $\pm$ 0.000 0.063 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 1.96 $\pm$ 0.00 5.32 $\pm$ 0.01 2.29 $\pm$ 0.00 0.060 $\pm$ 0.000 0.041 $\pm$ 0.000 0.058 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 1.92 $\pm$ 0.00 5.46 $\pm$ 0.01 2.26 $\pm$ 0.01 0.059 $\pm$ 0.000 0.039 $\pm$ 0.000 0.053 $\pm$ 0.000
Table 12: Set size and CovGap for CLIP (ViT-B) on 11 datasets across several few-shot adaptation approaches. Training for 4 shots.
@@
Aircraft ZS 18.70 $\pm$ 0.06 19.75 $\pm$ 0.06 20.74 $\pm$ 0.05 0.138 $\pm$ 0.000 0.139 $\pm$ 0.000 0.127 $\pm$ 0.000
ZSLP 9.73 $\pm$ 0.03 11.13 $\pm$ 0.03 11.55 $\pm$ 0.05 0.082 $\pm$ 0.001 0.073 $\pm$ 0.000 0.086 $\pm$ 0.001
CoOp (Zhou et al., 2022b) 9.41 $\pm$ 0.03 10.62 $\pm$ 0.04 11.38 $\pm$ 0.04 0.081 $\pm$ 0.001 0.076 $\pm$ 0.000 0.088 $\pm$ 0.001
KgCoOp (Hantao Yao, 2023) 9.46 $\pm$ 0.03 10.82 $\pm$ 0.03 11.00 $\pm$ 0.04 0.084 $\pm$ 0.001 0.076 $\pm$ 0.000 0.086 $\pm$ 0.001
CLAP (Silva-Rodríguez et al., 2024) 9.73 $\pm$ 0.03 11.00 $\pm$ 0.03 11.51 $\pm$ 0.05 0.082 $\pm$ 0.001 0.073 $\pm$ 0.000 0.086 $\pm$ 0.001
Caltech-101 ZS 0.94 $\pm$ 0.00 1.95 $\pm$ 0.01 1.20 $\pm$ 0.00 0.136 $\pm$ 0.001 0.092 $\pm$ 0.001 0.092 $\pm$ 0.001
ZSLP 0.91 $\pm$ 0.00 1.20 $\pm$ 0.00 1.05 $\pm$ 0.00 0.147 $\pm$ 0.001 0.085 $\pm$ 0.001 0.085 $\pm$ 0.001
CoOp (Zhou et al., 2022b) 0.92 $\pm$ 0.00 1.14 $\pm$ 0.00 1.06 $\pm$ 0.00 0.143 $\pm$ 0.001 0.085 $\pm$ 0.001 0.087 $\pm$ 0.001
KgCoOp (Hantao Yao, 2023) 0.92 $\pm$ 0.00 1.25 $\pm$ 0.00 1.08 $\pm$ 0.00 0.147 $\pm$ 0.001 0.084 $\pm$ 0.001 0.085 $\pm$ 0.001
CLAP (Silva-Rodríguez et al., 2024) 0.91 $\pm$ 0.00 1.31 $\pm$ 0.01 1.08 $\pm$ 0.00 0.144 $\pm$ 0.001 0.085 $\pm$ 0.001 0.086 $\pm$ 0.001
DTD ZS 11.78 $\pm$ 0.06 13.30 $\pm$ 0.04 13.66 $\pm$ 0.06 0.122 $\pm$ 0.001 0.117 $\pm$ 0.001 0.121 $\pm$ 0.001
ZSLP 4.74 $\pm$ 0.03 6.83 $\pm$ 0.03 5.67 $\pm$ 0.04 0.092 $\pm$ 0.001 0.077 $\pm$ 0.001 0.091 $\pm$ 0.001
CoOp (Zhou et al., 2022b) 5.68 $\pm$ 0.03 6.46 $\pm$ 0.04 7.00 $\pm$ 0.05 0.099 $\pm$ 0.001 0.093 $\pm$ 0.001 0.098 $\pm$ 0.001
KgCoOp (Hantao Yao, 2023) 5.33 $\pm$ 0.03 6.22 $\pm$ 0.04 6.33 $\pm$ 0.05 0.097 $\pm$ 0.001 0.083 $\pm$ 0.001 0.096 $\pm$ 0.001
CLAP (Silva-Rodríguez et al., 2024) 4.75 $\pm$ 0.03 6.83 $\pm$ 0.03 5.68 $\pm$ 0.04 0.094 $\pm$ 0.001 0.079 $\pm$ 0.001 0.093 $\pm$ 0.001
EuroSAT ZS 4.55 $\pm$ 0.01 4.78 $\pm$ 0.00 5.06 $\pm$ 0.01 0.103 $\pm$ 0.001 0.101 $\pm$ 0.001 0.102 $\pm$ 0.000
ZSLP 1.62 $\pm$ 0.00 2.22 $\pm$ 0.00 1.99 $\pm$ 0.00 0.043 $\pm$ 0.000 0.051 $\pm$ 0.000 0.042 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 1.82 $\pm$ 0.00 2.16 $\pm$ 0.00 1.95 $\pm$ 0.00 0.043 $\pm$ 0.000 0.028 $\pm$ 0.000 0.034 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 1.51 $\pm$ 0.00 1.92 $\pm$ 0.00 1.76 $\pm$ 0.00 0.028 $\pm$ 0.000 0.030 $\pm$ 0.000 0.032 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 1.73 $\pm$ 0.00 2.33 $\pm$ 0.00 2.06 $\pm$ 0.00 0.043 $\pm$ 0.000 0.048 $\pm$ 0.000 0.041 $\pm$ 0.000
Flowers-102 ZS 9.26 $\pm$ 0.05 9.52 $\pm$ 0.06 15.15 $\pm$ 0.08 0.180 $\pm$ 0.000 0.175 $\pm$ 0.000 0.179 $\pm$ 0.000
ZSLP 1.09 $\pm$ 0.00 2.36 $\pm$ 0.01 1.65 $\pm$ 0.00 0.132 $\pm$ 0.001 0.089 $\pm$ 0.001 0.095 $\pm$ 0.001
CoOp (Zhou et al., 2022b) 1.13 $\pm$ 0.00 1.80 $\pm$ 0.01 1.51 $\pm$ 0.00 0.104 $\pm$ 0.001 0.080 $\pm$ 0.001 0.083 $\pm$ 0.001
KgCoOp (Hantao Yao, 2023) 1.19 $\pm$ 0.00 2.04 $\pm$ 0.01 1.64 $\pm$ 0.00 0.114 $\pm$ 0.001 0.082 $\pm$ 0.001 0.085 $\pm$ 0.001
CLAP (Silva-Rodríguez et al., 2024) 1.15 $\pm$ 0.00 2.47 $\pm$ 0.01 1.70 $\pm$ 0.00 0.137 $\pm$ 0.001 0.092 $\pm$ 0.001 0.097 $\pm$ 0.001
Food-101 ZS 1.14 $\pm$ 0.00 1.89 $\pm$ 0.00 1.45 $\pm$ 0.00 0.054 $\pm$ 0.000 0.026 $\pm$ 0.000 0.031 $\pm$ 0.000
ZSLP 1.11 $\pm$ 0.00 1.82 $\pm$ 0.00 1.42 $\pm$ 0.00 0.046 $\pm$ 0.000 0.023 $\pm$ 0.000 0.026 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 1.15 $\pm$ 0.00 1.77 $\pm$ 0.00 1.41 $\pm$ 0.00 0.050 $\pm$ 0.000 0.026 $\pm$ 0.000 0.030 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 1.10 $\pm$ 0.00 1.86 $\pm$ 0.00 1.43 $\pm$ 0.00 0.050 $\pm$ 0.000 0.025 $\pm$ 0.000 0.027 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 1.10 $\pm$ 0.00 1.83 $\pm$ 0.00 1.42 $\pm$ 0.00 0.047 $\pm$ 0.000 0.023 $\pm$ 0.000 0.026 $\pm$ 0.000
ImageNet ZS 2.81 $\pm$ 0.00 10.07 $\pm$ 0.02 3.23 $\pm$ 0.00 0.086 $\pm$ 0.000 0.069 $\pm$ 0.000 0.084 $\pm$ 0.000
ZSLP 2.48 $\pm$ 0.00 8.02 $\pm$ 0.02 2.88 $\pm$ 0.00 0.078 $\pm$ 0.000 0.062 $\pm$ 0.000 0.078 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 2.69 $\pm$ 0.00 7.33 $\pm$ 0.01 3.03 $\pm$ 0.00 0.081 $\pm$ 0.000 0.067 $\pm$ 0.000 0.081 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 2.65 $\pm$ 0.00 8.60 $\pm$ 0.02 3.01 $\pm$ 0.00 0.082 $\pm$ 0.000 0.066 $\pm$ 0.000 0.082 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 2.44 $\pm$ 0.00 8.05 $\pm$ 0.01 2.82 $\pm$ 0.00 0.079 $\pm$ 0.000 0.062 $\pm$ 0.000 0.079 $\pm$ 0.000
Oxford Pets ZS 1.05 $\pm$ 0.00 1.43 $\pm$ 0.00 1.33 $\pm$ 0.00 0.109 $\pm$ 0.000 0.067 $\pm$ 0.000 0.072 $\pm$ 0.000
ZSLP 0.95 $\pm$ 0.00 1.26 $\pm$ 0.00 1.19 $\pm$ 0.00 0.080 $\pm$ 0.001 0.039 $\pm$ 0.000 0.038 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 0.95 $\pm$ 0.00 1.17 $\pm$ 0.00 1.13 $\pm$ 0.00 0.084 $\pm$ 0.000 0.036 $\pm$ 0.001 0.036 $\pm$ 0.001
KgCoOp (Hantao Yao, 2023) 0.94 $\pm$ 0.00 1.20 $\pm$ 0.00 1.15 $\pm$ 0.00 0.079 $\pm$ 0.000 0.038 $\pm$ 0.001 0.037 $\pm$ 0.001
CLAP (Silva-Rodríguez et al., 2024) 0.95 $\pm$ 0.00 1.28 $\pm$ 0.00 1.20 $\pm$ 0.00 0.084 $\pm$ 0.001 0.039 $\pm$ 0.000 0.038 $\pm$ 0.000
Cars ZS 2.24 $\pm$ 0.00 3.09 $\pm$ 0.01 2.49 $\pm$ 0.00 0.109 $\pm$ 0.000 0.080 $\pm$ 0.000 0.102 $\pm$ 0.000
ZSLP 1.62 $\pm$ 0.00 2.39 $\pm$ 0.01 1.94 $\pm$ 0.00 0.083 $\pm$ 0.000 0.061 $\pm$ 0.000 0.067 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 1.95 $\pm$ 0.00 2.64 $\pm$ 0.00 2.21 $\pm$ 0.00 0.087 $\pm$ 0.000 0.065 $\pm$ 0.000 0.078 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 1.99 $\pm$ 0.00 2.77 $\pm$ 0.00 2.29 $\pm$ 0.00 0.096 $\pm$ 0.000 0.069 $\pm$ 0.000 0.084 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 1.63 $\pm$ 0.00 2.40 $\pm$ 0.00 1.96 $\pm$ 0.00 0.085 $\pm$ 0.000 0.061 $\pm$ 0.000 0.067 $\pm$ 0.000
SUN397 ZS 2.83 $\pm$ 0.01 5.80 $\pm$ 0.01 3.21 $\pm$ 0.01 0.093 $\pm$ 0.000 0.073 $\pm$ 0.000 0.093 $\pm$ 0.000
ZSLP 2.00 $\pm$ 0.00 3.96 $\pm$ 0.01 2.29 $\pm$ 0.00 0.080 $\pm$ 0.000 0.060 $\pm$ 0.000 0.072 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 2.35 $\pm$ 0.00 3.76 $\pm$ 0.01 2.64 $\pm$ 0.00 0.082 $\pm$ 0.000 0.068 $\pm$ 0.000 0.081 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 2.34 $\pm$ 0.00 4.18 $\pm$ 0.01 2.60 $\pm$ 0.00 0.083 $\pm$ 0.000 0.065 $\pm$ 0.000 0.082 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 2.01 $\pm$ 0.00 3.99 $\pm$ 0.01 2.31 $\pm$ 0.00 0.080 $\pm$ 0.000 0.060 $\pm$ 0.000 0.072 $\pm$ 0.000
UCF101 ZS 2.89 $\pm$ 0.01 5.24 $\pm$ 0.02 3.83 $\pm$ 0.01 0.124 $\pm$ 0.000 0.096 $\pm$ 0.000 0.124 $\pm$ 0.001
ZSLP 1.58 $\pm$ 0.00 3.12 $\pm$ 0.01 1.98 $\pm$ 0.00 0.112 $\pm$ 0.000 0.074 $\pm$ 0.000 0.088 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 1.97 $\pm$ 0.01 3.12 $\pm$ 0.01 2.18 $\pm$ 0.01 0.114 $\pm$ 0.001 0.077 $\pm$ 0.000 0.110 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 1.74 $\pm$ 0.01 3.06 $\pm$ 0.01 2.06 $\pm$ 0.00 0.114 $\pm$ 0.001 0.070 $\pm$ 0.000 0.096 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 1.63 $\pm$ 0.01 3.25 $\pm$ 0.01 2.00 $\pm$ 0.00 0.114 $\pm$ 0.001 0.076 $\pm$ 0.000 0.090 $\pm$ 0.000